# A Single-Loop First-Order Algorithm for Linearly Constrained Bilevel Optimization

### v1

@@ Appendix A Useful Lemmas
Lemma A.1 (Lemma 12 in [3]).
@@
Lemma A.3 (Theorem 4.1 in [50]).
@@ Appendix C Reformulation
Similar results and proofs of Theorem˜4.1 can also be found in [21, 16]. For completeness, we also provide our proofs for Theorem˜4.1 here.
@@
The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [43]. According to Lemmas 2 and 3 in [16]. We know that when ˜3.1, 3.2, 3.3 hold, $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, defining $\psi(x)=g(x,y^{*}(x))$ and $\psi_{\delta}(x)=g_{\delta}(x,y^{*}_{\delta}(x))$, we have
@@
According to Theorem 2 in [45], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
@@
Since $g$ is strongly convex, $h$ is convex, $\nabla^{2}_{yy}g(x,y^{*}(x))\succ 0$, $\nabla^{2}_{yy}[\bar{h}(x,y^{*}(x))]_{i}\succeq 0$. Moreover, since $\bar{\lambda}^{*}(x)>0$, we have $\nabla^{2}_{yy}g(x,y^{*}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\succ 0$. Additionally, $\nabla_{y}\bar{h}(x,y^{*}(x))$ has full row rank. Thus, according to Lemma A.2 in [20], $H$ is invertible. We can calculate $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ with the following equation.
@@
According to [9], we have
@@ E.1 Potential function
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [16], we can find $\hat{v}$, $\hat{z}$ such that
@@ Appendix F Detailed Experimental Settings and Additional Experiments
We adapt and modify the code from [16]. The experiments on the toy example and hyperparameter optimization for SVM are conducted on an AMD EPYC 9554 64-Core Processor. The experiments on transportation network design are conducted on an Intel(R) Xeon(R) Platinum 8375C CPU. For the toy example, we set the hyperparameters for our algorithm as $\delta=0.1,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.01,\rho_{1}=\rho_{2}=1$.
@@ F.1 Hyperparameter optimization in SVM
We consider the same problem formulation as in [16], which is
@@
We compared our algorithm SFLCB with GAM [45], LV-HBA [49], BLOCC [16], and BiC-GAFFA [48] on the diabetes dataset. For GAM, we follow the same implementation approach and hyperparameters as [16], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [45]. For BLOCC and SFLCB, we set the LL objective as $\frac{1}{2}\|w\|^{2}+\frac{\mu}{2}b^{2}$, where $\mu=0.01$ serves as a regularization term to make the LL problem strongly convex w.r.t $w,b$. For LV-HBA, we set $\alpha=0.01,\gamma_{1}=0.1,\gamma_{2}=0.1,\eta=0.001$. For BLOCC, we set $\gamma=12,\eta=0.01,T=20,T_{y}=20,\eta_{1g}=0.001,\eta_{1F}=0.00001,\eta_{2g}=0.0001,\eta_{2F}=0.0001$. For BiC-GAFFA, we set $p=0.3,\gamma_{1}=10,\gamma_{2}=0.01,\eta=0.01,\alpha=0.001,\beta=0.001,c_{0}=10,R=10$. For our algorithm SFLCB, we set $\delta=0.01,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.001,\rho_{1}=\rho_{2}=0.01$. The hyperparameters of LV-HBA, BLOCC are the same as those used in [16] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [48]. The experiments are conducted across 10 different random train-validation-test splits, and the average results along with one standard deviation are reported in Figure˜2.
@@ F.2 Transportation network design
We consider the same setting as in [16]. In this setting, we act as the operator to design a new network that connects a set of stations $S$. Passengers then decide whether to use this network based on their rational decisions (lower level). The objective is to maximize the operator’s benefit (upper level). The operator can select a set of potential links $\mathcal{A}\subseteq S\times S$, and for each link $(i,j)\in\mathcal{A}$, determine its capacity $x_{ij}$. A link is constructed if $x_{ij}>0$; a larger $x_{ij}$ attracts more travelers, generating more revenue but incurring higher construction costs $c_{ij}$. Passenger demand is defined over a set of origin-destination pairs $\mathcal{K}\subseteq S\times S$. For each $(o,d)\in\mathcal{K}$, there is a known traffic demand $w_{od}$ and existing travel times $t^{\rm ext}_{od}$. We assume a single existing network. The fraction of passengers choosing the new network for each $(o,d)$ pair is denoted by $y_{od}$, and $y^{ij}_{od}$ represents the proportion using link $(i,j)$.
We summarize the notation as follows. We keep the notation the same as in [16].
@@
We consider the same 3 nodes and 9 nodes settings as in [16] and use the same lower level feasibility criteria. The hyperparameter settings used in Figure˜3 are listed below. For the 3 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.1$, $\rho_{1}=\rho_{2}=1000$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-4$, and $T=30000$. For the 9 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.25$, $\rho_{1}=\rho_{2}=50$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-5$, and $T=300000$. For BLOCC, we used the same hyperparameters as those used in [16].

### v2

@@ Appendix A Useful Lemmas
Lemma A.1 (Lemma 12 in [chen2021communication]).
@@
Lemma A.3 (Theorem 4.1 in [zhang2022global]).
@@ Appendix C Reformulation
Similar results and proofs of Theorem˜4.1 can also be found in [kwon2023penalty, jiang2024primal]. For completeness, we also provide our proofs for Theorem˜4.1 here.
@@
The uniqueness of Lagrangian multipliers is a direct result from the LICQ condition [wachsmuth2013licq]. According to Lemmas 2 and 3 in [jiang2024primal]. We know that when ˜3.1, 3.2, 3.3 hold, $\delta\leq\mu_{g}/(2l_{f,1})$, if, for all $x\in\mathcal{X}$, LICQ condition (Definition˜4.3) holds for $y^{*}(x)$ and $y^{*}_{\delta}(x)$, defining $\psi(x)=g(x,y^{*}(x))$ and $\psi_{\delta}(x)=g_{\delta}(x,y^{*}_{\delta}(x))$, we have
@@
According to Theorem 2 in [xu2023efficient], we know that when ˜3.1, 3.2, 3.3, 4.6 hold, if, for a given $x$, ˜4.5 and the LICQ (Definition˜4.3) condition holds for $y^{*}(x)$, then $\nabla\Phi(x)$ exists at $x$.
@@
Since $g$ is strongly convex, $h$ is convex, $\nabla^{2}_{yy}g(x,y^{*}(x))\succ 0$, $\nabla^{2}_{yy}[\bar{h}(x,y^{*}(x))]_{i}\succeq 0$. Moreover, since $\bar{\lambda}^{*}(x)>0$, we have $\nabla^{2}_{yy}g(x,y^{*}(x))+\nabla^{2}_{yy}\bar{h}(x,y^{*}(x))^{\top}\bar{\lambda}^{*}(x)\succ 0$. Additionally, $\nabla_{y}\bar{h}(x,y^{*}(x))$ has full row rank. Thus, according to Lemma A.2 in [kornowski2024first], $H$ is invertible. We can calculate $\nabla y^{*}(x)$ and $\lambda^{*}(x)$ with the following equation.
@@
According to [ghadimi2018approximation], we have
@@ E.1 Potential function
When h(x,y)=Bx+Aybh(x,y)=Bx+Ay-b, AA is full row rank and under ˜3.1, 3.2, 3.3 and $\delta=O(\epsilon)\leq\mu_{g}/(2l_{f,1})$ if we apply PGD for $\max_{v\in\mathbb{R}_{+}}\min_{z\in\mathbb{R}^{d_{y}}}g(x_{0},y)+v^{\top}(Bx_{0}+Az-b)$ with a fixed $x_{0}$, then according to Theorem 6 in [jiang2024primal], we can find $\hat{v}$, $\hat{z}$ such that
@@ Appendix F Detailed Experimental Settings and Additional Experiments
We adapt and modify the code from [jiang2024primal]. The experiments on the toy example and hyperparameter optimization for SVM are conducted on an AMD EPYC 9554 64-Core Processor. The experiments on transportation network design are conducted on an Intel(R) Xeon(R) Platinum 8375C CPU. For the toy example, we set the hyperparameters for our algorithm as $\delta=0.1,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.01,\rho_{1}=\rho_{2}=1$.
@@ F.1 Hyperparameter optimization in SVM
We consider the same problem formulation as in [jiang2024primal], which is
@@
We compared our algorithm SFLCB with GAM [xu2023efficient], LV-HBA [yao2024constrained], BLOCC [jiang2024primal], and BiC-GAFFA [yao2024overcoming] on the diabetes dataset. For GAM, we follow the same implementation approach and hyperparameters as [jiang2024primal], setting $\alpha=0.05,\epsilon=0.005$ and using a different formulation as introduced in [xu2023efficient]. For BLOCC and SFLCB, we set the LL objective as $\frac{1}{2}\|w\|^{2}+\frac{\mu}{2}b^{2}$, where $\mu=0.01$ serves as a regularization term to make the LL problem strongly convex w.r.t $w,b$. For LV-HBA, we set $\alpha=0.01,\gamma_{1}=0.1,\gamma_{2}=0.1,\eta=0.001$. For BLOCC, we set $\gamma=12,\eta=0.01,T=20,T_{y}=20,\eta_{1g}=0.001,\eta_{1F}=0.00001,\eta_{2g}=0.0001,\eta_{2F}=0.0001$. For BiC-GAFFA, we set $p=0.3,\gamma_{1}=10,\gamma_{2}=0.01,\eta=0.01,\alpha=0.001,\beta=0.001,c_{0}=10,R=10$. For our algorithm SFLCB, we set $\delta=0.01,\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=0.001,\rho_{1}=\rho_{2}=0.01$. The hyperparameters of LV-HBA, BLOCC are the same as those used in [jiang2024primal] and the hyperparameters of BiC-GAFFA are the same as those used in their original paper [yao2024overcoming]. The experiments are conducted across 10 different random train-validation-test splits, and the average results along with one standard deviation are reported in Figure˜2.
@@ F.2 Transportation network design
We consider the same setting as in [jiang2024primal]. In this setting, we act as the operator to design a new network that connects a set of stations $S$. Passengers then decide whether to use this network based on their rational decisions (lower level). The objective is to maximize the operator’s benefit (upper level). The operator can select a set of potential links $\mathcal{A}\subseteq S\times S$, and for each link $(i,j)\in\mathcal{A}$, determine its capacity $x_{ij}$. A link is constructed if $x_{ij}>0$; a larger $x_{ij}$ attracts more travelers, generating more revenue but incurring higher construction costs $c_{ij}$. Passenger demand is defined over a set of origin-destination pairs $\mathcal{K}\subseteq S\times S$. For each $(o,d)\in\mathcal{K}$, there is a known traffic demand $w_{od}$ and existing travel times $t^{\rm ext}_{od}$. We assume a single existing network. The fraction of passengers choosing the new network for each $(o,d)$ pair is denoted by $y_{od}$, and $y^{ij}_{od}$ represents the proportion using link $(i,j)$.
We summarize the notation as follows. We keep the notation the same as in [jiang2024primal].
@@
We consider the same 3 nodes and 9 nodes settings as in [jiang2024primal] and use the same lower level feasibility criteria. The hyperparameter settings used in Figure˜3 are listed below. For the 3 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.1$, $\rho_{1}=\rho_{2}=1000$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-4$, and $T=30000$. For the 9 nodes setting, we set the hyperparameters of our method SFLCB as: $\delta=0.25$, $\rho_{1}=\rho_{2}=50$, $\eta_{x}=\eta_{y}=\eta_{z}=\eta_{u}=\eta_{v}=3e-5$, and $T=300000$. For BLOCC, we used the same hyperparameters as those used in [jiang2024primal].